# Deep Neural OFDM Receivers: Two Novel Architectures for BER-BLER Optimization and Comparison with State-of-the-Art Architectures

### v2

@@ Abstract
Neural receivers have recently become a popular topic, where the received signals can be directly decoded by data-driven mechanisms such as machine learning and deep learning. In this paper, we propose two novel neural network–based orthogonal frequency division multiplexing (OFDM) receivers performing channel estimation and equalization tasks and directly predicting log-likelihood ratios (LLRs) from the received in-phase and quadrature-phase (IQ) signals. The first network, the Dual Attention Transformer (DAT), employs a state-of-the-art (SOTA) transformer architecture with an attention mechanism. The second network, the Residual Dual Non-Local Attention Network (RDNLA), utilizes a parallel residual architecture with a non-local attention block. The bit-error rate (BER) and block-error rate (BLER) performance of various SOTA neural receiver architectures is compared with our proposed methods across different signal-to-noise ratio (SNR) levels. The simulation results show that DAT and RDNLA outperform both traditional communication systems and existing neural receiver models. The computational efficiency of the proposed neural receivers supports their feasibility for next-generation communication systems.
@@
Data-driven techniques such as machine learning (ML) and deep learning (DL) have seen significant advances recently, not only in foundational algorithms but also in hardware design perspective, mainly driven by the “artificial intelligence (AI) at the edge” concept. For example, tensor accelerators along with new AI algorithms make computationally challenging tasks feasible to be deployed in real-world applications. These advancements facilitate an increasing number of AI deployements at the edge [1]. Wireless communication community has also been heavily investing AI/ML based applications, where they can either build new foundational models or adopt existing ones from other domains by training them using wireless communication datasets.
Considering advanced radio access network (RAN) technologies along with the usage of larger bandwidths in the upcoming 3GPP releases, signal processing requirements of enormous data in real-time make a traditional mathematical-based system design with cascaded communication blocks consisting heuristic algorithms prohibitively expensive. This brings an AI-native system design forefront, where the AI/ML approaches can focus on the most important features of the underlying system by appropriately setting the weights of the trained models which provide a direct mapping between transmitted (input) and received (output) information. In addition, the AI/ML models are suitable for revealing highly complex features for several advanced RAN functions, where hardware accelerators play a critical role in processing enormous data in parallel [2]. As long as high-quality data is available, the model is trained and deployed at the edge. The 3GPP has already started to study AI/ML adaptation for wireless systems [3].
@@
The recent studies on a tensor native wireless communication [7] and ray tracing based simulation framework Sionna [8] provide important advancements in the context of neural receivers. The potential of neural receiver architectures has been demonstrated in [9] and [10], where neural receivers are utilized in 5G multi-user multiple input-multiple output (MIMO) scenarios in real-time while adhering to the standard compliance. Deep learning-based channel estimation methods [11], incorporating belief knowledge and transfer learning strategies [12] for single input-multiple output (SIMO) receivers [13] further improve performance by adapting to realistic channel conditions, as verified by aerial experiments [13]. In addition, interpretability analysis [14] and performance verification frameworks for AI-specific transceiver actions [15] provide important guidance for the robustness, safety, and operational reliability of these models. Also, [16] showed neural receiver performance can be further improved by adopting new DL models such as transformers.
@@
  • We introduce two deep learning models for neural receivers that significantly improve BER and block-error rate (BLER) performance compared to the conventional Transformers. Proposed models also demonstrate notable improvements over other state-of-the-art (SOTA) models.
@@
  • Lastly, we will make publicly available the source code of our proposed models together with other SOTA model implementations [^1] https://github.com/Hisar-Research/sionna-nn-receiver-playground. This allows researchers and practitioners to explore, test, and expand upon our findings.
@@
I System Model and Background
@@
From that point, after receiving OFDM RG, the goal of this study is to estimate the LLR values of received constellation points with DL techniques. However, before proceeding with DL model, we will give traditional receiver blocks required to estimate LLR values. The UE transmits pilot symbols at specific subcarriers and time slots, forming a pilot pattern in the OFDM RG $\mathbf{X}_{p}$. The received pilot signals at the BS are
@@
and the least squares (LS) channel estimate at the pilot locations is given by,
@@
Figure 2: The architectures of proposed LLR estimation methods for neural receiver.
II Proposed DL-Based LLR Estimation Methods
@@
II-A Input Preprocessing and Embedding
@@
II-B Dual Attention Transformer
@@
II-B1 Multi-Head Self Attention
@@
$[\mathbf{Q},\,\mathbf{K},\,\mathbf{V}]$ $=\mathbf{Z}\,\mathbf{U}_{\mathrm{QKV}}\;\in\;\mathbb{R}^{T\times 3D}.$
@@
II-B2 Dual-Attention Mechanism
@@
Finally, the last dense layer unit size is set to $N_{\text{bits}}$ in the last FC layer to match the desired shape $\mathbf{L}$, where the output is reshaped into $\mathbf{Z_{out}}\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times 2}$. In short, DAT emphasizes learning token-wise (spatial) and feature-wise (channel) correlations via self-attention and dual attention blocks. It is particularly adept at capturing long-range dependencies across subcarriers and symbols.
II-C Residual Dual Non-Local Attention Network
@@
II-C1 Parallel Residual Block
@@
II-C2 Dual Non-Local Attention Block
@@
where $\mathcal{S}(\mathbf{X})$ and $\mathcal{C}(\mathbf{X})$ denote the restored-channel outputs of the spatial and channel branches, respectively. This dual-path design allows each location to attend to all others across both spatial and channel dimensions, while the skip connection preserves identity information. The network’s output can be obtained by adding the output of the first convolutional layer to the output of each block and then passing through the last convolutional layer, as shown below:
$\mathbf{Z}_{\text{out }}=\operatorname{Conv_{1}}\left(\mathbf{F}_{\mathrm{PRB}_{2}}+\mathbf{F}_{0}\right)\ \in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times 2}.$ (32)
@@
II-D Model Training
The given models are trained using the bit-metric decoding (BMD) rate, which serves as an information theoretic objective for bit-interleaved coded modulation (BCIM) systems. The BMD rate is computed from the transmitted coded bits and the predicted LLRs, and its optimization ensures accurate soft-bit estimation for subsequent decoding. Binary-cross entropy (BCE) loss function compares the predicted LLRs with the true transmitted coded bits over the entire OFDM RG, which is given by
@@ II-D Model Training
Figure 3: BER (a) and BLER (b) of the proposed neural receivers with state-of-the art models at varying noise levels.
III Simulation Experiments and Results
@@
This section presents the simulation experiments and results using Nvidia’s open-source Sionna library. A wireless communication system consisting of a single antenna UE ($N_{\mathrm{UE}}=1$) and a BS with two antennas $N_{\mathrm{BS}}=2$ in Fig. 1 is generated within the Sionna simulator using a single stream configuration. The OFDM system consists of 76 subcarriers and 14 OFDM symbols per each frame, where 2 symbols are allocated for pilots. A cyclic prefix of 6 samples is used to mitigate inter-symbol interference. The signal is modulated using QAM, while LDPC encoding and decoding are applied at a coding rate of $R_{C}=0.5$. The proposed two neural receiver architectures, namely DAT and RDNLA, are implemented within the Sionna simulator in addition to eight existing SOTA DL architectures from the literature. BER and BLER results for ten architectures are shown as $E_{b}/N_{0}$ is swept from -3 dB to 5 dB during simulations, however the plots confined to specific interval for clarity. Each performance result corresponds to the average of $3200$ Monte-Carlo experiments, with the target block error run set to 1000. For training, the batch size is set to 128 while the number of epochs is $10^{5}$.
@@
Fig. 3(a) shows the BER results of ten different receiver models including our proposed DAT and RDNLA models under different $E_{b}/N_{0}$ values. The figure distinguishes two groups of receiver architectures. The first group (ResNet [6], Transformer [16], ResDNLA, DAT) achieves the BER results below $10^{-4}$ even $E_{b}/N_{0}$ is lower than 0.5 dB. This success is due to their architectural components such as local, global attention blocks, and residual connections, which effectively capture complex and global relationships and enhance feature representation for neural receiver tasks. The second group (EffNetV2, NAFNet ResUNet, RIDNet[18, 19, 20, 21]) have demonstrated lower BER compared to their counterpart, especially EffNet-V2, NAFNet and ResUNet. Although these models perform well in semantic segmentation and incorporate skip connections and conventional attention mechanisms, they underperform in LLR estimation due to their inability to effectively capture both local and global dependencies required for this task. The first group excels at modelling both time-frequency and spatial–channel relationships, while DAT and RDNLA further outperform standard Transformer and ResNet architectures by using finely crafted attention mechanisms that effectively capture complex feature dependencies and preserve global features. DAT and RDNLA are able to achieve $10^{-4}$ BER below 0 dB while Transformer and Resnet can achieve this result above this point. This can be seen as a minor gain however trend shows that in higher SNR regimes DAT and RDNLA draw away especially after 0 dB because Transformer and ResNet are not able to learn local dependencies at these points.
@@
TABLE I: Mean inference time of models per OFDM frame
@@
Table I shows that all models process an OFDM frame within a narrow window (±1.65 ms), reflecting a balanced design in layer depth and attention complexity. The Transformer is the fastest, as it relies on dense layers and self-attention, both dominated by efficient matrix multiplications. In contrast, convolutional models (ResNet, ResUNet, RIDNet, RDNLA) use many small kernel operations that are memory-bound.
IV Discussion and Future Work
This paper proposes two neural network–based OFDM receivers and compares their BER and BLER performance with eight existing state-of-art models, where the DAT employs a transformer architecture with an attention mechanism while the RDNLA utilizes a parallel residual architecture with a non-local attention block. The simulation experiments are conducted using the open-source Sionna simulator. The results show that the proposed models outperform both traditional receiver systems and existing neural receivers. Also, results show that BER alone is not enough to train the model because it ignores how errors cluster, a factor that heavily impacts BLER. Models that disperse bit errors tend to achieve lower BLER even with higher BER because scattered errors are easier for the decoder to correct, whereas clustered errors can overwhelm it and trigger entire block failures. Thus, we will analyze the spatial and temporal patterns of decoding errors and design architectures and training schemes that promote error dispersion. In particular, we plan to incorporate error-aware loss functions and decoder-in-the-loop feedback to directly optimize for BLER.

### v3

# Novel Deep Neural OFDM Receiver Architectures for LLR Estimation

@@ Abstract
Neural receivers have recently become a popular topic, where the received signals can be directly decoded by data-driven mechanisms such as machine learning and deep learning. In this paper, we propose two novel neural network–based orthogonal frequency division multiplexing (OFDM) receivers performing channel estimation and equalization tasks and directly predicting log-likelihood ratios (LLRs) from the received in-phase and quadrature-phase (IQ) signals. The first network, the Dual Attention Transformer (DAT), employs a state-of-the-art (SOTA) transformer architecture with an attention mechanism. The second network, the Residual Dual Non-Local Attention Network (RDNLA), utilizes a parallel residual architecture with a non-local attention block. The bit-error rate (BER) and block-error rate (BLER) performance of various SOTA neural receiver architectures is compared with our proposed methods across different signal-to-noise ratio (SNR) levels. The simulation results show that DAT and RDNLA outperform both traditional communication systems and existing neural receiver models.
@@
I Introduction
Data-driven techniques such as machine learning (ML) and deep learning (DL) have seen significant advances recently, not only in foundational algorithms but also in hardware design perspective, mainly driven by the “artificial intelligence (AI) at the edge” concept. For example, tensor accelerators along with new AI algorithms make computationally challenging tasks feasible to be deployed in real-world applications. These advancements facilitate an increasing number of AI deployments at the edge [1]. The wireless communication community has also been heavily investing in AI/ML-based applications, where they can either build new foundational models or adapt existing ones from other domains by training them using wireless communication datasets.
Considering advanced radio access network (RAN) technologies along with the usage of larger bandwidths in the upcoming 3GPP releases, signal processing requirements of enormous data in real-time make a traditional mathematical-based system design with cascaded communication blocks consisting of heuristic algorithms prohibitively expensive. This brings an AI-native system design forefront, where the AI/ML approaches can focus on the most important features of the underlying system by appropriately setting the weights of the trained models which provide a direct mapping between transmitted (input) and received (output) information. In addition, the AI/ML models are suitable for revealing highly complex features for several advanced RAN functions, where hardware accelerators play a critical role in processing enormous data in parallel [2]. As long as high-quality data is available, the model is trained and deployed at the edge. The 3GPP has already started to study AI/ML adaptation for wireless systems [3].
@@ I Introduction
The recent studies on a tensor native wireless communication [7] and Nvidia’s ray tracing based simulation framework Sionna [8] provide important advancements in the context of neural receivers. The potential of neural receiver architectures has been demonstrated in [9] and [10], where neural receivers are utilized in 5G multi-user multiple input-multiple output (MIMO) scenarios in real-time while adhering to the standard compliance. Deep learning-based channel estimation methods [11], incorporating belief knowledge and transfer learning strategies [12] for single input-multiple output (SIMO) receivers [13] further improve performance by adapting to realistic channel conditions, as verified by aerial experiments [13]. In addition, interpretability analysis [14] and performance verification frameworks for AI-specific transceiver actions [15] provide important guidance for the robustness, safety, and operational reliability of these models. Also, [16] showed neural receiver performance can be further improved by adopting new DL models such as transformers.
@@
  • We introduce two deep learning models for neural receivers that significantly improve BER and block-error rate (BLER) performance compared to the conventional Transformers. Proposed models also demonstrate notable improvements over other state-of-the-art (SOTA) models. Additionally, we provided the computation times of each model with the conventional least square (LS) estimation receiver.
@@
  • Lastly, we will make the source code of our proposed models together with other SOTA model implementations publicly available [^1] https://github.com/Hisar-Research/sionna-nn-receiver-playground. This allows researchers and practitioners to explore, test, and expand upon our findings.
@@
II System Model and Background
@@
From that point, after receiving OFDM RG, the goal of this study is to estimate the LLR values of received constellation points with DL techniques. However, before proceeding with the DL model, we will give traditional receiver blocks required to estimate LLR values. The UE transmits pilot symbols at specific subcarriers and time slots, forming a pilot pattern in the OFDM RG $\mathbf{X}_{p}$. The received pilot signals at the BS are
@@
and the LS channel estimate at the pilot locations is given by,
@@
Figure 2: The DAT (left) and RDNLA (right) LLR estimation architectures for neural receiver.
III Proposed DL-Based LLR Estimation Methods
@@
III-A Input Preprocessing and Embedding
@@
III-B Dual Attention Transformer
@@
III-B1 Multi-Head Self Attention
@@
$[\mathbf{Q},\,\mathbf{K},\,\mathbf{V}]$ $=\mathbf{Z}\,\mathbf{U}_{\mathrm{QKV}}\;\in\;\mathbb{R}^{N_{T}\times 3N_{D}}.$
@@
III-B2 Dual-Attention Mechanism
@@
Finally, the last dense layer unit size is set to $N_{\text{bits}}$ in the last FC layer to match the desired shape $\mathbf{L}$, where the output is reshaped into $\mathbf{Z_{out}}\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times N_{\text{bits}}}$. In short, DAT emphasizes learning token-wise (spatial) and feature-wise (channel) correlations via self-attention and dual attention blocks. It is particularly adept at capturing long-range dependencies across subcarriers and symbols.
III-C Residual Dual Non-Local Attention Network
@@
III-C1 Parallel Residual Block
@@
III-C2 Dual Non-Local Attention Block
@@
where $\mathcal{S}(\mathbf{X})$ and $\mathcal{C}(\mathbf{X})$ denote the restored-channel outputs of the spatial and channel branches, respectively. This dual-path design allows each location to attend to all others across both spatial and channel dimensions, while the skip connection preserves identity information. The network’s output can be obtained by adding the output of the first convolutional layer to the output of each block and then passing through the last convolutional layer with $N_{\text{bits}}$ filters, as shown below
$\mathbf{Z}_{\text{out }}=\operatorname{Conv_{1}}\left(\mathbf{F}_{\mathrm{PRB}_{2}}+\mathbf{F}_{0}\right)\ \in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times N_{\text{bits}}}.$ (32)
@@
III-D Model Training
The given models are trained using the bit-metric decoding (BMD) rate, which serves as an information theoretic objective for bit-interleaved coded modulation (BCIM) systems. The BMD rate is computed from the transmitted coded bits and the predicted LLRs, and its optimization ensures accurate soft-bit estimation for subsequent decoding. The binary-cross entropy (BCE) loss function compares the predicted LLRs with the true transmitted coded bits over the entire OFDM RG, which is given by
@@ III-D Model Training
Figure 3: BER (a) and BLER (b) of the proposed neural receivers with SOTA models at varying noise levels.
IV Simulation Experiments and Results
@@
This section presents the simulation experiments and results using Nvidia’s open-source Sionna library. A wireless communication system consisting of a single antenna UE ($N_{\mathrm{UE}}=1$) and a BS with two antennas $N_{\mathrm{BS}}=2$ in Fig. 1 is generated using a single stream configuration. The OFDM system consists of 76 subcarriers and 14 OFDM symbols per frame, where 2 symbols are allocated for pilots. A cyclic prefix of 6 samples is used to mitigate inter-symbol interference. The signal is modulated using QAM, while LDPC encoding and decoding are applied at a coding rate of $R_{C}=0.5$. The proposed two neural receiver architectures, namely DAT and RDNLA, are implemented in addition to eight existing SOTA DL architectures from the literature. BER and BLER for all ten architectures are plotted as functions of $E_{b}/N_{0}$; for clarity, the BER curves span from -2.5 dB to 1.5 dB, while the BLER curves cover -1.5 dB to 2 dB. Each performance result corresponds to the average of $3200$ Monte-Carlo experiments, with the target block error run set to 1000. For training, the batch size is set to 128 while the number of epochs is $10^{5}$.
@@
Fig. 3(a) shows the BER results of ten different receiver models including our proposed DAT and RDNLA models under different $E_{b}/N_{0}$ values. Two separate performance clusters of receiver architectures emerge in the figure. The first group (ResNet [6], Transformer [16], ResDNLA, DAT) achieves the BER results below $10^{-4}$ even $E_{b}/N_{0}$ is lower than 0.5 dB. This success is due to their architectural components such as local, and global attention blocks, and residual connections, which effectively capture complex and global relationships and enhance feature representation for neural receiver tasks. The second group (EffNet-V2, NAFNet ResUNet, RIDNet [18, 19, 20, 21]) have demonstrated lower BER compared to their counterpart, especially EffNet-V2, NAFNet and ResUNet. Although these models perform well in semantic segmentation and incorporate skip connections and conventional attention mechanisms, they underperform in LLR estimation due to their inability to effectively capture both local and global dependencies required for this task. The first group excels at modelling both time-frequency and spatial–channel relationships, while DAT and RDNLA further outperform standard Transformer and ResNet architectures by using finely crafted attention mechanisms that effectively capture complex feature dependencies and preserve global features. DAT and RDNLA achieve a BER of $10^{-4}$ at $E_{b}/N_{0}$ below 0 dB. In contrast, the Transformer and ResNet require $E_{b}/N_{0}$ above 0 dB to reach the same. While the gain is modest at low $E_{b}/N_{0}$ (below -1 dB), beyond 0 dB DAT and RDNLA steadily pull ahead, as the Transformer and ResNet fall into decline in capturing local dependencies at higher SNR.
@@
TABLE I: Mean inference times per OFDM RG
@@
LS-Estimation 0.14
Table I shows that all models process an OFDM RG within a narrow window (±1.65 ms), reflecting a balanced design in layer depth and attention complexity. The Transformer is the fastest, as it relies on dense layers and self-attention, both dominated by efficient matrix multiplications. In contrast, convolutional models (ResNet, ResUNet, RIDNet, RDNLA) use many small kernel operations that are memory-bound. Not surprisingly, the conventional LS-estimation receiver is faster than the neural models, underscoring the trade-off between computational efficiency and performance.
V Discussion and Future Work
This paper proposes two neural network–based OFDM receivers and compares their BER and BLER performance with eight existing state-of-art models, where the DAT employs a transformer architecture with an attention mechanism while the RDNLA utilizes a parallel residual architecture with a non-local attention block. The results show that the proposed models outperform both traditional receiver systems and existing neural receivers. However, neural receivers still require efficient architectures to catch the speed of conventional receivers. In addition, the results show that BER alone is insufficient to train the model because it ignores how errors cluster, a factor that heavily impacts BLER. The models that disperse bit errors tend to achieve lower BLER even with higher BER because scattered errors are easier for the decoder to correct, whereas the clustered errors can overwhelm it and trigger entire block failures. Thus, we will analyze the spatial and temporal patterns of decoding errors and design architectures and training schemes that promote error dispersion. In particular, we plan to incorporate error-aware loss functions and decoder-in-the-loop feedback to directly optimize BLER.